OpenStudy (anonymous):

express the perimeter of the triangle as a polynomial side 1:) 2x-3 side 2:)x-12 side 3:)7x+1

jimthompson5910 (jim_thompson5910):

perimeter = sum of all the sides so you just add up all the sides...ie the 3 expressions

OpenStudy (anonymous):

so would it be... 9x-14

jimthompson5910 (jim_thompson5910):

well focus on the x terms

jimthompson5910 (jim_thompson5910):

add up 2x, x and 7x to get _____

OpenStudy (anonymous):

9x^3?

jimthompson5910 (jim_thompson5910):

2x+x is 3x 3x+7x = 10x

jimthompson5910 (jim_thompson5910):

so.... 2x+x+7x is 10x

jimthompson5910 (jim_thompson5910):

you do NOT say x+x is x^2 or x+x+x = x^3 that's not correct

OpenStudy (anonymous):

ok I for got that just x is the same as 1x. so the answer is 10x-14

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

forgot*

jimthompson5910 (jim_thompson5910):

just remember that saying x+x is like saying " i have one thing...then I add another thing....so I now have two things total"

OpenStudy (anonymous):

also what determines a polynomial? ex. which expression is a polynomial 1) \[\frac{ 3 }{\sqrt{x} }\] 2) \[\frac{ 2 }{ x }+3x\] 3)\[\sqrt{x} -6\] 4)\[x-\sqrt{6}\]

jimthompson5910 (jim_thompson5910):

if you're dividing by a variable, then it's NOT a polynomial if the exponent over the variable isn't a whole number, then it's NOT a polynomial

jimthompson5910 (jim_thompson5910):

if you have square roots over the variable, then it's NOT a polynomial

OpenStudy (anonymous):

so #2 is a polynomial?

jimthompson5910 (jim_thompson5910):

are you dividing by a variable in #2 ?

jimthompson5910 (jim_thompson5910):

ie...is a variable in the denominator

OpenStudy (anonymous):

so not 1 or 2 but 4?

jimthompson5910 (jim_thompson5910):

yep, there is a square root...but it's not over the variable

OpenStudy (anonymous):

OK